# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v2

@@ III.1 Tight-binding $s-d$ model for antiferromagnet and altermagnet
Figure 6: Density profiles for the split system $\varepsilon_{2(3)}=\pm\delta$ showing Friedel oscillations at $\mu=-1.5t$. Both the unit cell-resolved density (a) and row-resolved density (d) show clear Friedel oscillations, but the row-resolved density also show the $\pi$-density wave due to the presence of non-zero on-site potentials. The Fourier transform of the unit cell-resolved density is shown in (b) and the row-resolved in (e). Several peaks arise in both corresponding to the dominant frequencies in the density oscillations. We label the dominant peaks with the notation $2k_{F,i}$ where i is a simple index. Due to the 2D geometry, we do not have one particular, but rather a distribution of $2k_{F}^{x}$ from states from different y-modes. We quantify this distribution by summing the row-resolved bulk spectral function evaluated at the Fermi level across all y-momenta, shown in (f) where $k_{x}$ runs over the full Brillouin zone (FBZ), obtaining a $2k_{F}^{x}$ distribution which matches well with the observed frequency content in the density oscillations. Note that values of $2k_{F}^{x}$ beyond $\pi/2$ appear in the reduced Brillouin zone (RBZ) in (c) due to aliasing and is distinguished by an asterisk ($2k_{F}^{*}$). (e) In the row-resolved density, an additional set of frequencies given by $|2k_{F}\pm\pi|$ arise due to the fact that the staggered $\pi$ order is modulated in magnitude by the underlying eigenfunctions.
@@
Consider now a normal metal (NM) lattice with $J_{\text{sd}}=0$. We consider two distinct configurations of the non-magnetic on-site potentials in order to elucidate the interplay between the unit cell potentials and the lattice geometry. In Fig. 5, we plot the unit cell-resolved and row-resolved densities for a system with no on-site potentials in (a) and (b) and a split potential $\varepsilon_{2(3)}=\pm\delta$ in (c) and (d) where $\delta/t\ll 1$ (see the insets for unit cell layout). The densities are summed over the two first standing-wave harmonics in x for the homogeneous y-mode ($k=0$). The first observation is that the densities on sites in the left and right column of the unit cell are shifted by a lattice constant $a$ [see Fig. 5(a) and (c)]. In a unit cell picture where both columns correspond to the same indices $(I,k)$, this shift originating from the relative x-displacement between columns, manifests as a magnitude difference between the density profiles. The second observation is that the magnitudes of the $\alpha$-resolved densities are altered individually by the respective on-site potentials. In the case of $\varepsilon_{2}=\varepsilon_{3}=0$, all sites in the unit cell are equivalent causing the only distinguishing factor to be the one lattice constant $a$ shift between densities in different columns. For the split configuration ($\varepsilon_{2(3)}=\pm\delta$), this equivalence is lifted. In Fig. 5(c), the $\alpha=2$ density is suppressed as $\varepsilon_{2}=\delta$ constitute an effective lowering of the chemical potential. Along the same lines, an enhancement is observed for $\alpha=3$ due to $\varepsilon_{3}=-\delta$. The magnitudes of densities at sites $\alpha=1,4$ sites also change due to a changed unit cell environment, but the effect is equal across the two sites and comparatively smaller. To understand the significance of these relative shifts in magnitude, the row-resolved picture is illustrative. The row-resolved densities are obtained by plotting the densities of unit cell sites in the same row ($\alpha=1,2$ and $\alpha=3,4$) together, shifted relatively by a lattice constant $a$, shown in Figs. 5(b) and 5(d). As is evident from Fig. 5(b), the unit cell description of the NM with $\varepsilon_{2}=\varepsilon_{3}=0$ does not essentially provide any new information not present in the original model. The advantage of the unit-cell picture arise when on-site potentials are introduced. While the unit cell-resolved densities (Fig. 5(c)) are modulated in magnitude due to the on-site potentials, they retain their smooth standing wave-like profiles as all sites of the same $\alpha$ see the same potential change. The row-resolved density, however, is given by alternating left and right column sites, and the row-resolved densities in Fig. 5(d) are now distinctly different, where a $\pi$-periodic density wave emerge on each row due to the alternating magnitude modulation. When the sum in Eq. (35) is extended to a large number of eigenfunctions, the row-resolved density profiles can become intricate. As we have seen, however, the observed oscillations can be deconvoluted in the unit cell picture as the combination of a one lattice constant $a$ shift and magnitude differences between densities in different columns, caused by the chosen on-site potential configuration and unit cell geometry.
@@
In a system with broken translation invariance the oscillations of the charge density known as Friedel oscillations [55, 56] arise in the vicinity of interfaces and defects. While the oscillations of eigenfunctions interfere and cancel out in the bulk of the ribbon, the open boundaries force all eigenfunctions to be roughly in phase close to the boundaries leading to the Friedel oscillations. The frequency of 1D Friedel oscillations is $2k_{F}$ where $k_{F}$ is the Fermi momentum. In 2D, the Fermi surface does in general give rise to a distribution of Fermi wave vectors, and the whole distribution of $k_{F}^{x}$, the x component of the Fermi wave vectors, can leave a mark on the density oscillations.
@@
To exemplify how these density oscillations occur, the density for the split potential ($\varepsilon_{2(3)}=\pm\delta$) is shown in Fig. 6 for the unit-cell- 6(a) and row-resolved 6(d) representations. Both show clear oscillations in the vicinity of the interface while roughly homogeneous in the ribbon bulk. In Figs. 6(b) and 6(e), the Fourier transform of the density in the two representations is shown, indicating the dominant frequencies in the density oscillations. These frequencies correspond to roughly twice the x-component of the Fermi vector for the homogeneous y-mode. Due to the 2D Fermi surface, however, eigenfunctions with a range of $k_{F}^{x}$ exist at the Fermi surface, giving rise to additional peaks in the frequency spectrum. These can be shown to correspond closely to the distribution of $k_{F}^{x}$ in the partially integrated spectral function evaluated at the Fermi surface, obtained by integrating over all y-momenta, show in Figs. 6(c) and 6(f). Note in particular that the presence of the split potentials break the symmetry between the rows. As such, the row-resolved Fermi-surfaces are identical in shape, but with a slight difference in the distribution of states on the Fermi surface. This is evident in Fig. 6(c) and 6(f) [see insets in 6(f)] where the distributions of x-components are different on the two rows.
@@
Finally, we point out that the density oscillations arising from the Fermi surface are largely independent on the specific on-site potentials as long as these are small relative to t. As long as the Fermi surface does not change significantly upon changing the potentials, the system eigenfunctions will exhibit similar Friedel oscillations, but with an additional staggered order on top. The non-zero potentials leave a more direct mark on the row-resolved density oscillations [Fig. 6(e)] through the emergence of a $\pi$-peak, due to the previously discussed mechanism, and through the appearance of additional peaks at frequencies $|2k_{F}\pm\pi|$, caused by the $2k_{F}$ modulation of the $\pi$ density wave magnitude. In effect, the presence of periodic potentials on the lattice gives rise to a magnitude modulation on top of the standing wave eigenfunctions. The magnitude of these modulations follow the magnitude of the eigenfunction as a whole, causing a coupling between the underlying $2k_{F}$ frequency of the initial eigenfunction, and the $\pi$-frequency arising from the lattice geometry. This gives rise to the additional peaks in the Fourier transform of the row-resolved density in Fig. 6(d) close to $k_{x}=0$.
The purpose of this section is to identify the key mechanisms at play in the microscopic model concerning density modulations. With an understanding of the mechanisms that determine the oscillation pattern of the electron density near an interface, we are now ready to move on to discuss the consequences of spin-dependent potentials in the unit cell.
@@ III.3 Non-zero $J_{\text{sd}}$ coupling and AFM order
Figure 7: Spin-up and spin-down densities are shown in (a) for an AFM with $J_{\text{sd}}=t$, $\mu=-2.75t$. The spin-dependent $J_{\text{sd}}$ gives rise to highly spin-dependent occupations on the magnetic sites $\alpha=1,4$. (b) The magnitude differences in the underlying eigenfunctions gives rise to a staggered magnetization (green and orange). The magnetization is out of phase between the two rows causing the total unit cell magnetization (black) to approach zero in the ribbon center. Despite no spin-dependence in the Friedel oscillations, we still observe uncompensated magnetization (black) in the vicinity of the interface. This is caused by an effective coupling between the staggered order and the spin-independent Friedel oscillations, not captured in the continuum model. Note that we here plot the magnitude of the unit cell magnetization multiplied by a factor 5 for clarity.
For a given spin-species, a non-zero $J_{\text{sd}}$ has a very similar effect as the split potentials discussed above and gives rise to a $\pi$-periodic density wave. The s-d coupling is, however, spin-dependent and changes sign between the two spin species. Due to the unit cell geometry, the spin-up and spin-down density waves are out of phase, giving rise to a staggered magnetization persisting through the bulk of the ribbon. In our model, this is the mechanism responsible for the AFM staggered order in the itinerant electrons, a microscopic character which is absent in the effective model in Sec. II.
@@
In Fig. 7(a), we plot the spin-up and spin-down densities. In Fig. 7(b), we plot the staggered (row-resolved) and unit cell magnetizations, where the latter is averaged over all unit cell sites. While the unit cell magnetization practically vanishes in the center of the ribbon, significant deviations from the compensated bulk order arise in the vicinity of the interfaces. As the effective continuum model of AFM considered in Sec. II.1 does not capture the uncompensated nature of the interface nor the staggered order, it omits this type of coupling between Friedel oscillations and magnetization.
We point out that the Fermi momentum itself is spin-independent in the AFM and one thus do not expect any magnetization contribution from Fermi surface mismatch, which the continuum model confirms. The frequency of the density and magnetization oscillations are shown in detail in Fig. 9. Notice in particular that we observe the unit cell magnetization to oscillate at $2k_{F}$ [9(d)] exactly due to the effective coupling between the $\pi$-order and Friedel oscillations. Therefore, one of the ways to achieve a surface-induced magnetization in continuum models would be to introduce spin-dependent boundary conditions, which would correspond to a different environment of spin-up and spin-down sites in the lattice model.
Following this reasoning, the emergence of uncompensated edge magnetization for itinerant electrons in an AFM depend critically on the unit cell geometry and the nature of the interface. The interface-induced magnetization arise only if the up/down spins interact differently with the interface, creating a spin-dependent environment which does not vanish upon averaging over the interface. To achieve this, an uncompensated interface is typically needed. For a compensated interface, there will still be relative shifts between the spin-up and down density waves, but the shifts will vanish when summing over a unit cell as both spins interact with the interface symmetrically. Along these lines, we also expect this effect to be vulnerable towards surface structure and roughness. As the magnetization is induced solely through the one lattice constant shift, a roughening of the surface could cancel out the spin-dependent interface environment. This is discussed in detail in Sec. III.6.
@@ III.4 Phenomenological model for magnetization oscillations in antiferromagnets
Motivated by the results in the lattice model, where the spin-up and spin-down sites experience different environment and are shifted with respect to each other by a lattice constant, we propose a phenomenological model. We use the model and the results of Sec. II.1, however, we shift the coordinates for the spin-down electrons by a lattice constant $a$. Then the magnetization is determined by
@@
While not being rigorous, the model captures the shift between the spin-up and spin-down populations crucial for the lattice model. Furthermore, one can view the shift by a lattice constant as modified boundary conditions where the wave functions of spin-up and spin-down electrons acquire different values at the interfaces.
For wide ribbons, $Lk_{F}\gg 1$, the magnetization away from the edges $xk_{F}\gg 1$ reads
$m_{z}(x)\!$ =$\stackrel{{\scriptstyle\tilde{x}\to\infty}}{{=}}$ $\!\frac{m_{0}}{\sqrt{\pi}}\Bigg{[}\frac{\cos{\left(2\tilde{x}+\frac{\pi}{4}\right)}}{\tilde{x}^{3/2}}-\frac{\cos{\left(2\tilde{x}+2\tilde{a}+\frac{\pi}{4}\right)}}{(\tilde{x}+\tilde{a})^{3/2}}\Bigg{]}\!+\mathcal{O}{\left(\frac{1}{\tilde{x}^{2}}\right)}$ (37)
=$\stackrel{{\scriptstyle\tilde{a}\to 0}}{{=}}$ $\!\frac{2m_{0}\tilde{a}}{\sqrt{\pi}}\frac{1}{\tilde{x}^{3/2}}\sin{\left(2\tilde{x}+\frac{\pi}{4}\right)}+\mathcal{O}{\left(\tilde{a}^{2},\frac{1}{\tilde{x}^{2}}\right)}.$
The oscillations of the magnetization in the phenomenological model of antiferromagnets are (i) suppressed as $\tilde{a}=ak_{F}\ll 1$, (ii) shifted by $\pi/2$ in phase with respect to the particle density oscillations, and (iii) decay as $1/\tilde{x}^{3/2}$ at $\tilde{x}\to\infty$. In addition, the phenomenological model predicts vanishing net magnetization in antiferromagnets.

### v3

@@ III.1 Tight-binding s-d model for antiferromagnet and altermagnet
Figure 6: The density profiles for the split system $\varepsilon_{2(3)}=\pm\delta$ showing Friedel oscillations at $\mu=-1.5t$. Both the unit-cell-resolved density (a) and row-resolved density (d) show clear Friedel oscillations, but the row-resolved density also shows the $\pi$-density wave due to the presence of non-zero on-site potentials. The Fourier transform of the unit-cell-resolved density is shown in (b) and the row-resolved in (e). Several peaks arise in both corresponding to the dominant frequencies in the density oscillations. We label the dominant peaks with the notation $2k_{F,i}$ where i is a simple index. Due to the 2D geometry, we do not have one particular, but rather a distribution of $2k_{F}^{x}$ from states from different y-modes. We quantify this distribution by summing the row-resolved bulk spectral function evaluated at the Fermi level across all y-momenta, shown in (f) where $k_{x}$ runs over the full Brillouin zone (FBZ), obtaining a $2k_{F}^{x}$ distribution which matches well with the observed frequency content in the density oscillations. Note that values of $2k_{F}^{x}$ beyond $\pi/2$ appear in the reduced Brillouin zone (RBZ) in (c) due to aliasing and are distinguished by an asterisk ($2k_{F}^{*}$). (e) In the row-resolved density, an additional set of frequencies given by $|2k_{F}\pm\pi|$ arise due to the fact that the staggered $\pi$ order is modulated in magnitude by the underlying eigenfunctions.
@@
Consider now a normal metal (NM) lattice with $J_{\text{sd}}=0$. We examine two distinct configurations of the nonmagnetic on-site potentials in order to elucidate the interplay between the unit-cell potentials and the lattice geometry. In Fig. 5, we plot the unit-cell-resolved and row-resolved densities for a system with no on-site potentials in (a) and (b) and a split potential $\varepsilon_{2(3)}=\pm\delta$ in (c) and (d) where $\delta/t\ll 1$ (see the insets for unit-cell layout). The densities are summed over the two first standing-wave harmonics in x for the homogeneous y-mode ($k=0$). The first observation is that the densities on sites in the left and right column of the unit-cell are shifted by a lattice constant $a$ [see Fig. 5(a) and (c)]. In a unit-cell picture where both columns correspond to the same indices $(I,k)$, this shift originating from the relative x-displacement between columns, manifests as a magnitude difference between the density profiles. The second observation is that the magnitudes of the $\alpha$-resolved densities are altered individually by the respective on-site potentials. In the case of $\varepsilon_{2}=\varepsilon_{3}=0$, all sites in the unit-cell are equivalent causing the only distinguishing factor to be the one lattice constant $a$ shift between densities in different columns. For the split configuration ($\varepsilon_{2(3)}=\pm\delta$), this equivalence is lifted. In Fig. 5(c), the $\alpha=2$ density is suppressed as $\varepsilon_{2}=\delta$ represents an effective lowering of the chemical potential. Along the same lines, an enhancement is observed for $\alpha=3$ due to $\varepsilon_{3}=-\delta$. The magnitudes of densities at sites $\alpha=1,4$ sites also change due to a changed unit-cell environment, but the effect is equal across the two sites and comparatively smaller. To understand the significance of these relative shifts in magnitude, the row-resolved picture is illustrative. The row-resolved densities are obtained by plotting the densities of unit-cell sites in the same row ($\alpha=1,2$ and $\alpha=3,4$) together, shifted relatively by a lattice constant $a$, shown in Figs. 5(b) and 5(d). As is evident from Fig. 5(b), the unit-cell description of the NM with $\varepsilon_{2}=\varepsilon_{3}=0$ does not essentially provide any new information not present in the original model. The advantage of the unit-cell picture arises when on-site potentials are introduced. While the unit-cell-resolved densities (Fig. 5(c)) are modulated in magnitude due to the on-site potentials, they retain their smooth standing wave-like profiles as all sites of the same $\alpha$ see the same potential change. The row-resolved density, however, is given by alternating left and right column sites, and the row-resolved densities in Fig. 5(d) are now distinctly different, where a $\pi$-periodic density wave emerges on each row due to the alternating magnitude modulation. When the sum in Eq. (36) is extended to a large number of eigenfunctions, the row-resolved density profiles can become intricate. As we have seen, however, the observed oscillations can be deconvoluted in the unit-cell picture as the combination of a one lattice constant $a$ shift and magnitude differences between densities in different columns, caused by the chosen on-site potential configuration and unit-cell geometry.
@@
In a system with broken translation invariance, the oscillations of the charge density known as Friedel oscillations [55, 56] arise in the vicinity of interfaces and defects. While the oscillations of eigenfunctions interfere and cancel out in the bulk of the ribbon, the open boundaries force all eigenfunctions to be roughly in phase close to the boundaries leading to the Friedel oscillations. The frequency of 1D Friedel oscillations is $2k_{F}$ where $k_{F}$ is the Fermi momentum. In 2D, the Fermi surface does in general give rise to a distribution of Fermi wave vectors, and the whole distribution of $k_{F}^{x}$, the x component of the Fermi wave vectors, can leave a mark on the density oscillations.
@@
To exemplify how these density oscillations occur, the density for the split potential ($\varepsilon_{2(3)}=\pm\delta$) is shown in Fig. 6 for the unit-cell- 6(a) and row-resolved 6(d) representations. Both show clear oscillations in the vicinity of the interface while roughly homogeneous in the ribbon bulk. In Figs. 6(b) and 6(e), the Fourier transform of the density in the two representations is shown, indicating the dominant frequencies in the density oscillations. These frequencies correspond to roughly twice the x-component of the Fermi vector for the homogeneous y-mode. Due to the 2D Fermi surface, however, eigenfunctions with a range of $k_{F}^{x}$ exist at the Fermi surface, giving rise to additional peaks in the frequency spectrum. These can be shown to correspond closely to the distribution of $k_{F}^{x}$ in the partially integrated spectral function evaluated at the Fermi surface, obtained by integrating over all y-momenta, shown in Figs. 6(c) and 6(f). Note in particular that the presence of the split potentials breaks the symmetry between the rows. As such, the row-resolved Fermi surfaces are identical in shape, but with a slight difference in the distribution of states on the Fermi surface. This is evident in Fig. 6(c) and 6(f) [see insets in 6(f)] where the distributions of x-components are different on the two rows.
@@
Finally, we point out that the density oscillations arising from the Fermi surface are largely independent of the specific on-site potentials as long as these are small relative to $t$. As long as the Fermi surface does not change significantly upon changing the potentials, the system eigenfunctions will exhibit similar Friedel oscillations, but with an additional staggered order on top. The non-zero potentials leave a more direct mark on the row-resolved density oscillations [Fig. 6(e)] through the emergence of a $\pi$ peak, due to the previously discussed mechanism, and through the appearance of additional peaks at frequencies $|2k_{F}\pm\pi|$, caused by the $2k_{F}$ modulation of the $\pi$-density wave magnitude. In effect, the presence of periodic potentials on the lattice gives rise to a magnitude modulation on top of the standing wave eigenfunctions. The magnitude of these modulations follows the magnitude of the eigenfunction as a whole, causing a coupling between the underlying $2k_{F}$ frequency of the initial eigenfunction, and the $\pi$ frequency arising from the lattice geometry. This gives rise to the additional peaks in the Fourier transform of the row-resolved density in Fig. 6(d) close to $k_{x}=0$.
@@ III.3 Non-zero $J_{\text{sd}}$ coupling and AFM order
Figure 7: The spin-up and spin-down densities are shown in (a) for an AFM with $J_{\text{sd}}=t$, $\mu=-2.75t$. The spin-dependent $J_{\text{sd}}$ gives rise to highly spin-dependent occupations on the magnetic sites $\alpha=1,4$. (b) The magnitude differences in the underlying eigenfunctions give rise to a staggered magnetization (green and orange). The magnetization is out of phase between the two rows causing the total unit-cell magnetization (black) to approach zero in the ribbon center. Despite no spin dependence in the Friedel oscillations, we still observe uncompensated magnetization (black) in the vicinity of the interface. This is caused by an effective coupling between the staggered order and the spin-independent Friedel oscillations, not captured in the continuum model. Note that we here plot the magnitude of the unit-cell magnetization multiplied by a factor of 5 for clarity. Inset shows a zoom-in on the first 20 sites at the left interface.
For a given spin projection, a non-zero $J_{\text{sd}}$ has a very similar effect as the split potentials discussed above and gives rise to a $\pi$-periodic density wave. The s-d coupling is, however, spin-dependent and changes the sign between the two spin species. Due to the unit-cell geometry, the spin-up and -down density waves are out of phase, giving rise to a staggered magnetization persisting through the bulk of the ribbon. In our model, this is the mechanism responsible for the AFM staggered order in the itinerant electrons, a microscopic character which is absent in the effective model in Sec. II.
@@
In Fig. 7(a), we plot the spin-up and spin-down densities. In Fig. 7(b), we plot the staggered (row-resolved) and unit-cell magnetizations, where the latter is averaged over all unit-cell sites. While the unit-cell magnetization practically vanishes in the center of the ribbon, significant deviations from the compensated bulk order arise in the vicinity of the interfaces. As the effective continuum model of AFM considered in Sec. II.1 does not capture the uncompensated nature of the interface nor the staggered order, it omits this type of coupling between Friedel oscillations and magnetization.
We point out that the Fermi momentum itself is spin-independent in the AFM and one thus does not expect any magnetization contribution from Fermi surface mismatch, which the continuum model confirms. The frequency of the density and magnetization oscillations are shown in detail in Fig. 8. Notice in particular that we observe the unit-cell magnetization to oscillate at $2k_{F}$ [8(d)] exactly due to the effective coupling between the $\pi$-order and Friedel oscillations. Therefore, one of the ways to achieve surface-induced magnetization in continuum models would be to introduce spin-dependent boundary conditions, which would correspond to a different environment of spin-up and spin-down sites in the lattice model.
Figure 8: (a) The spin-up and -down row-resolved densities for the AFM in Fig. 7 are Fourier transformed to obtain the frequencies of the density oscillations. The dominant frequency is denoted $2k_{F,1}$. By considering the spectral function integrated over all y momenta and evaluated at the Fermi level (c), the dominant frequency can be seen to correspond closely to the peak in the distribution of $k_{F}^{x}$, i.e., x components of the Fermi vector of states at the Fermi level. The additional peak in the spectrum in (a) arises because the magnitude of the $\pi$ staggered order itself is modulated by the density oscillations, giving rise to a $|2k_{F}\pm\pi|$ peak. (b) The row-resolved magnetization contains the same frequencies as those of the underlying spin-densities while the dominant frequency of the unit-cell magnetization (d) is $2k_{F,1}^{*}$, corresponding to the reduced Brillouin zone representation of $2k_{F,1}$, caused by aliasing.
Following this reasoning, the emergence of uncompensated edge magnetization for itinerant electrons in an AFM depends critically on the unit-cell geometry and the nature of the interface. The interface-induced magnetization arises only if the up and down spins interact differently with the interface, creating a spin-dependent environment which does not vanish upon averaging over the interface. To achieve this, an uncompensated interface is typically needed. For a compensated interface, there will still be relative shifts between the spin-up and -down density waves, but the shifts will vanish when summing over a unit-cell as both spins interact with the interface symmetrically. Along these lines, we also expect this effect to be vulnerable to surface structure and roughness. As the magnetization is induced solely through the one lattice constant shift, a roughening of the surface could cancel out the spin-dependent interface environment. This is discussed in detail in Sec. III.6.
@@ III.4 Phenomenological model for magnetization oscillations in antiferromagnets
Motivated by the results in the lattice model, where the spin-up and -down sites experience different environments and are shifted with respect to each other by a lattice constant, we propose a phenomenological model. We use the continuum model and the results of Sec. II.1, however, we shift the coordinates for the spin-down electrons by a lattice constant $a$. Then the magnetization is determined by
@@
While not being rigorous, the model captures the shift between the spin-up and -down populations crucial for the lattice model. Furthermore, one can view the shift by a lattice constant as modified boundary conditions where the wave functions of spin-up and -down electrons acquire different values at the interfaces.
For wide ribbons, $\tilde{L}\gg 1$, the magnetization away from the edges $\tilde{x}\gg 1$ reads as
$m_{z}(x)$ $=$ $\frac{m_{0}}{\sqrt{\pi}}\Bigg{[}\frac{\cos{\left(2\tilde{x}+\frac{\pi}{4}\right)}}{\tilde{x}^{3/2}}-\frac{\cos{\left(2\tilde{x}+2\tilde{a}+\frac{\pi}{4}\right)}}{(\tilde{x}+\tilde{a})^{3/2}}\Bigg{]}\!+\mathcal{O}{\left(\frac{1}{\tilde{x}^{2}}\right)}$ (38)
$=$ $\frac{2m_{0}\tilde{a}}{\sqrt{\pi}}\frac{1}{\tilde{x}^{3/2}}\sin{\left(2\tilde{x}+\frac{\pi}{4}\right)}+\mathcal{O}{\left(\tilde{a}^{2},\frac{1}{\tilde{x}^{2}}\right)},$
where we expanded in $\tilde{x}\gg 1$ in the first line and in $\tilde{a}\ll 1$ in the second line. The oscillations of the magnetization in the phenomenological model of antiferromagnets are (i) suppressed as $\tilde{a}=ak_{F}\ll 1$, (ii) shifted by $\pi/2$ in phase with respect to the particle density oscillations, and (iii) decay as $1/\tilde{x}^{3/2}$ at $\tilde{x}\to\infty$. In addition, the phenomenological model predicts vanishing net magnetization in antiferromagnets.